OpenStudy (anonymous):

THIS SHOULD BE SO SIMPLE FOR ME BUT ITS NOT MEDAL+FAN What is the solution to the system of equations? x+y=5 x-2y=2 (___), (___)

OpenStudy (mathwizzard3):

I think the answer is no solution.

OpenStudy (anonymous):

its not multiple choice it only has a place for me to type the solution. :/

OpenStudy (phi):

why don't you multiply the bottom equation by -1 (that means both sides and *every* term by -1) then add the two equations what do you get ?

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

i'm not sure...

OpenStudy (phi):

multiply the bottom equation by -1 what do you get ?

OpenStudy (anonymous):

-x-(-2y)=-2?

OpenStudy (phi):

yes, and -(-2y) is +2y, so you can write it as -x+2y= -2 now add the top equation: x+ y=5 -x+2y= -2 can you add them ?

OpenStudy (anonymous):

just add 5 and -2?

OpenStudy (phi):

yes, on the right side but also add the left sides and put an = sign in

OpenStudy (anonymous):

3y=3?

OpenStudy (phi):

yes because x+-x + y + 2y simplifies to 0+3y or 3y on the left side, and 5-2 =3 on the right side 3y=3 now divide both sides by 3

OpenStudy (anonymous):

1y=1?

OpenStudy (anonymous):

or y=1

OpenStudy (phi):

yes, but people usually don't bother to write 1y. they say y y=1 now use either equation (I choose the simplest one) x+y=5 put in 1 instead of y: x+1= 5 now add -1 to both sides and simplify

OpenStudy (anonymous):

x=4?

OpenStudy (phi):

yes, so the solution is x=4, y=1 or as an (x,y) pair, the answer is (4,1)

OpenStudy (anonymous):

thank you so much!! I have one more question like this one can you help me?

OpenStudy (phi):

if you type fast

OpenStudy (anonymous):

ok lol ill open another question